# FreshStack: Building Realistic Benchmarks for Evaluating Retrieval on Technical Documents

### v1

@@
Abstract.
We introduce FreshStack, a reusable framework for automatically building information retrieval (IR) evaluation benchmarks from community-asked questions and answers. FreshStack conducts the following steps: (1) automatic corpus collection from code and technical documentation, (2) nugget generation from community-asked questions and answers, and (3) nugget-level support, retrieving documents using a fusion of retrieval techniques and hybrid architectures. We use FreshStack to build five datasets on fast-growing, recent, and niche topics to ensure the tasks are sufficiently challenging. On FreshStack, existing retrieval models, when applied out-of-the-box, significantly underperform oracle approaches on all five topics, denoting plenty of headroom to improve IR quality. In addition, we identify cases where rerankers do not clearly improve first-stage retrieval accuracy (two out of five topics). We hope that FreshStack will facilitate future work toward constructing realistic, scalable, and uncontaminated IR and RAG evaluation benchmarks. FreshStack datasets are available at: https://fresh-stack.github.io.
Information Retrieval, General Framework, Evaluation Benchmark
1. Introduction
Retrieval-augmented generation (RAG) is a popular technique to enhance traditional information retrieval (IR) capabilities with language model generation. RAG systems use large language models (LLMs) to generate long-form responses (Guu et al., 2020; Lewis et al., 2020; Izacard and Grave, 2021; Borgeaud et al., 2022), grounded in the information available from retrieved documents (Khandelwal et al., 2020; Lewis et al., 2020; Gao et al., 2023b; Liu et al., 2024a). Despite its wide usage, evaluating RAG remains incredibly challenging. Existing IR and RAG benchmarks are not well-suited for evaluation, as these are outdated and highly limited. In particular, we observe three major issues in existing benchmarks:
  • Lack of realistic, open-ended questions: Existing datasets contain purely extractive short answers (e.g., Natural Questions (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017)) or crowd-sourced questions (e.g., HotPotQA (Yang et al., 2018)). A limited number of datasets capture “natural” human-asked questions, i.e., MS MARCO (Nguyen et al., 2016) or Natural Questions (Kwiatkowski et al., 2019), but unfortunately, brief and straightforward questions are inserted into a search box, failing to represent the complex questions that real users might pose to modern RAG systems.
@@ 1. Introduction
  • Static and unspecialized: After sourcing questions and answers, a benchmark becomes at the risk of (1) contamination, if current LLMs are trained on the same set of documents or questions, (2) overfitting, when systems inevitably saturate by repeated internal or external leaderboarding (e.g., BEIR (Thakur et al., 2021)), and (3) staleness, when questions or answers are not refreshed and become outdated.
@@
Figure 1. A data instance from LangChain generated with FreshStack. The question and answer pair is sourced from Stack Overflow and provided as input to generate nuggets with GPT-4o (highlighted in blue). Next, code snippets and technical documents from multiple GitHub repositories (e.g., Jupyter Notebook) are chunked, processed, and pooled for each question. Finally, each pooled document chunk is judged with GPT-4o for binary relevance (either yes or no) at a nugget-level, i.e., whether the document factually supports the information present in each nugget.
IR / RAG Benchmarks Niche Complex Dynamic Challenge
Domains Questions Updates Level
CQADupstack (Hoogeveen et al., 2015) No No No Easy
CodeSearchNet (Husain et al., 2019) No No No Easy
COIR (Li et al., 2024a) Limited Yes No Moderate
Stack Overflow-QA (Li et al., 2024a) No Yes No Moderate
CodeRAG-Bench (Wang et al., 2024a) Limited No No Moderate
Neural Code Search (Li et al., 2019) No No No Moderate
SWE-Bench (Jimenez et al., 2024) No Yes Yes High
FreshStack (ours) Yes Yes Yes High
Table 1. A comparison of existing IR/RAG evaluation benchmarks with FreshStack.
@@
A realistic benchmark should measure model generalization on niche domains, and continue to update. Additionally, it must capture the complexity of human-generated queries—such as multi-hop reasoning (Yang et al., 2018), code understanding (Jimenez et al., 2024), or specialized terminology—rather than relying on artificially easy questions. This drives the robustness of systems in answering questions in evolving public libraries or private code bases (Zhang et al., 2023a; Jimenez et al., 2024), a company’s internal forum (Raina and Gales, 2024), or technical troubleshooting (Soman and Roychowdhury, 2024; Pu et al., 2025; Chen et al., 2025).
@@
In our work, we introduce FreshStack, a holistic framework for constructing realistic datasets on niche and challenging domains, seeking to avoid contamination due to (perpetual) recency. Using FreshStack, we construct an evaluation benchmark on five niche topics sourced from community-asked questions and answers on Stack Overflow and a corpus containing code snippets and technical documents from public GitHub repositories. The FreshStack framework contains three major steps:
  1. (1)
    Automatic corpus collection (Section 3.2) with technical documents chunked and sourced from several GitHub repositories.
  2. (2)
    Nugget generation (Section 3.3) with GPT-4o using community-asked questions and answers in Stack Overflow.
  3. (3)
    Nugget-level support (Section 3.5) with GPT-4o on document chunks, retrieved from a fusion of retrieval techniques and hybrid architectures.
FreshStack has several advantages over existing benchmarks (as highlighted in Table 1). First, the framework utilizes user-asked questions and curated answers, making the evaluation challenging. We are not crafting artificial (LLM-generated) questions or sampling questions myopically. Second, all answers are supported in real time by information from technical documentation in GitHub repositories. Third, the framework is designed to be general and scalable without modification. Finally, the framework is focused on niche domains and recent topics, taking careful measures to mitigate risks with data contamination introduced by LLMs.
We provide a complete overview of our framework and investigate three research questions (RQs) to provide insights into the design choices used in FreshStack:
  • RQ1
    How to construct realistic and challenging IR evaluation datasets from user-asked questions?
  • RQ2
    How do LLMs act as an assessor for (1) nugget generation with community-asked questions & answers and (2) nugget-level support with GitHub-sourced documents?
  • RQ3
    How do out-of-the-box retrieval and reranker systems perform on benchmark generated with FreshStack?
@@
We investigate the framework’s quality by constructing five datasets on niche topics, e.g., LangChain or Godot. We calibrate the automatic stage with GPT-4o using a machine learning (ML) expert, assessing the quality of nugget generation and nugget-level support for one of the topics (LangChain). Our results show that LLM-generated nuggets capture crucial information required to answer the question, and GPT-4o precisely labels support at a nugget level.
@@
For the judgment pool construction, we assess pooling results by comparing the oracle (having access to the answer) and inference (relying only on the question) settings, showing that ensemble fusion in question decomposition (in the inference setting) and nugget generation (in the oracle setting) outperforms other techniques, respectively. Beyond pool construction, we explore the document retrieval setting on FreshStack, evaluating retrieval and rerankers out-of-the-box, using the question alone. We find that retrieval models drastically underperform oracle systems on all five topics, showing a high headroom for improvement on specialized topics. In addition, ensemble fusion outperforms individual models, indicating that diversity in models enhances retrieval, and rerankers provide clear benefits in some but not all topics.
FreshStack is a general framework and can be applied to any domain of a similar structure. The datasets will be made publicly available to the research community. Overall, we hope FreshStack can serve as a testbed for future work to develop similar challenging benchmarks for evaluating IR and RAG systems.
2. Related Work
@@
Retrieval-augmented generation. RAG has been widely used to avoid “hallucinations” (Zhang et al., 2023b) seen by LLMs when handling knowledge-intensive tasks (Kandpal et al., 2023). RAG reduces the generation of factually incorrect content, leading to adoption in various commercial systems, e.g., Bing Search or Google AI Overviews. Existing IR and RAG benchmarks are stale, focused on text, evaluating on academic question answering datasets (Gao et al., 2023b; Rosenthal et al., 2025; Ram et al., 2023), or are not challenging, being constructed for RAG (Yang et al., 2024; Chen et al., 2024; Li et al., 2024b; Niu et al., 2024). A limited number of datasets refresh over time to avoid LLM decontamination (Kasai et al., 2023; Vu et al., 2024; Shao et al., 2024; White et al., 2025), however, these contain easy and unrealistic questions for evaluation. In contrast, FreshStack generates niche and challenging datasets, which are not constructed for RAG, and can refresh over time.
@@
Code-based benchmarks. Neural code generation (Lu et al., 2021) requires LLMs to generate code from scratch for generic programming questions. One popular benchmark is SWE-Bench (Jimenez et al., 2024), which evaluates whether LLMs can generate code changes for GitHub pull requests (PRs) in popular public repositories. Similarly, CodeSearchNet (Husain et al., 2019), COIR (Li et al., 2024a), LiveCodeBench (Jain et al., 2024), and CodeRAG-Bench (Wang et al., 2024a) focus on the evaluation of high-level programming problems on popular public repositories. In contrast, in FreshStack, we focus on assisting developers, from a novice to a domain expert, by providing real-time answers on fast-growing and recent topics such as LangChain (introduced in 2023) by referencing technical documentation in GitHub repositories.
@@
Stack Overflow datasets. FreshStack is not the first dataset to use Stack Overflow for retrieval. However, the evaluation setting of retrieving canonical documents from GitHub repositories remains under-explored. Existing datasets such as CQADupstack (Hoogeveen et al., 2015), LoTTE (Santhanam et al., 2022), and Stack Overflow-QA (Li et al., 2024a) follow a different task, to retrieve the relevant answer snippet given the question asked by a real user on Stack Overflow. The closest setting similar to FreshStack is found in Neural Code Search (Li et al., 2019), which incorporates public documentation and code snippet examples from GitHub as the corpus to answer questions asked by real users on popular programming topics such as Android.
@@
Figure 2. The FreshStack framework: (1) Stack Overflow questions and answers are sourced for recent and niche topics. (2) GitHub repository documents are collected and chunked to form the corpus (for each topic). (3) Nuggets or atomic facts within the question and answer are generated with GPT-4o. (4) Ensemble techniques and models retrieve documents, which construct our document judgment pools. (5) GPT-4o evaluates support for every document-nugget pair as a binary judgment.
3. The FreshStack Framework
@@
The FreshStack framework involves five stages to construct an end-to-end evaluation dataset (highlighted in Figure 2). The framework includes three major design choices:
  1. (1)
    Inclusion of recent and niche topics actively discussed in computer programmer communities such as Stack Overflow.
  2. (2)
@@
    General and automatic framework that can be extended to different domains (potentially even across languages) without much manual effort.
  3. (3)
@@
Stack Overflow is an online question answering platform for computer programmers. It functions as a collaborative knowledge base, utilizing a voting system to curate content quality. Users ask questions about a particular topic and provide a description (often a code snippet with the error message) with a Stack Overflow tag. Anyone in the community can answer the question, however, higher-voted answers are prioritized, reflecting community consensus on the accuracy and relevance of the answer. Questions and answers are also tagged by topics, allowing for easy retrieval of topic-wise questions.
3.1. Stack Overflow Topic and Question Selection
@@
For topic selection, we target niche and recent topics introduced on Stack Overflow from 2023 onward, containing a minimum of 50 posts. We sort all topics using the overall number of posts and curate five topics starting from the highest (LangChain) to the lowest frequency (Yolo v7 & v8) covering different domains and sufficiently different from each other.
@@
Questions & Answers. We extract relevant posts and answers from the Stack Overflow XML data archive (dated October 2024). [^1] The Stack Overflow XML data archive (CC BY-SA license) is updated once every quarter: https://meta.stackexchange.com/questions/401324/announcing-a-change-to-the-data-dump-process We scan the archive to pick all the relevant posts as questions containing the required tag and to filter answer posts to the questions. At the end, we filter and keep questions with an accepted answer, prioritizing precision over quantity in retaining questions with high-quality answers.
3.2. Automatic Corpus Collection with GitHub
@@
Answering a higher percentage of questions requires a robust set of corpora from multiple sources. For instance, addressing issues in LangChain may require ChromaDB GitHub documentation to resolve errors related to its usage. In our work, we build a different corpus set per topic by combining multiple GitHub repositories as sources (we list the GitHub repositories per topic in Table 2).
Table 2. GitHub repositories used for constructing the document collection for every topic in FreshStack.
@@
Stack Overflow Tags. We analyze which GitHub repository to select for the document corpus by analyzing tag frequency from Stack Overflow posts. This involves identifying top-k co-occurring tags, where k is the threshold balancing question coverage with indexing costs. Some tags are generic, such as Python, whereas others are specific, such as LangChainJS. Filtering the tags to keep only a subset of repositories does not degrade the FreshStack dataset quality. We manually verify GitHub repositories for each tag, with plans to automate this procedure in the future. We could have constructed the corpus directly, however, the top-k co-occurring tags provide a useful signal on relevant GitHub repositories.
@@
Chunking & Indexing. We clone the latest branch of a GitHub repository in a local workspace and parse all files as a tree structure. Each file (either a text document or code snippet) is chunked into small sections of up to a maximum of 4096 tokens, skipping non-text formats. [^2] We skip indexing images, videos, .bin, .csv, and audio files or unrecognized file formats. The GitHub filepath serves as the document identifier, with additional chunk details encoded in the identifier. Finally, we combine all chunks into a single corpus, prefixing all document identifiers with the repository name to identify the common files in each repository separately (e.g., LICENSE or requirements.txt).
3.3. Nuggetization or Nugget Generation
@@
A nugget is a core concept or atomic fact essential in a system’s response. The term nugget was informally referred to as SCU (summary content units) as clauses appearing in model summarization (Nenkova and Passonneau, 2004) and later formalized as “information nugget” for evaluating long-form answers (Voorhees, 2003; Lin and Demner-Fushman, 2005, 2006; Pavlu et al., 2012). Nuggetization refers to constructing or generating nuggets from information-dense text. The procedure decomposes a verbose answer into key atomic facts or essential components, aiding evaluation. More recently, with the onset of RAG, nugget-based evaluation has renewed interest with LLMs for factual accuracy assessment in long answers (Arabzadeh and Clarke, 2024; Mayfield et al., 2024; Farzi and Dietz, 2024; Pradeep et al., 2024).
@@
Nuggetization. We automatically generate nuggets from Stack Overflow question-answer pairs using GPT-4o (OpenAI, 2024), avoiding the cumbersome procedure of manual nugget construction (Pradeep et al., 2024). LLM-based nugget generation is explored in the TREC 2024 RAG track [^3] TREC 2024 RAG track: https://trec-rag.github.io/ (Pradeep et al., 2024) and in multiple works (Dietz, 2024; Farzi and Dietz, 2024). Separately, we experimented with prompting techniques and found that grading notes style prompts (Liu et al., 2024b) provided parseable and high-quality nuggets in our experiments.
3.4. Retrieval: Oracle & Inference Setting

### v2

@@
Abstract
We introduce FreshStack, a holistic framework for automatically building information retrieval (IR) evaluation benchmarks by incorporating challenging questions and answers. FreshStack conducts the following steps: (1) automatic corpus collection from code and technical documentation, (2) nugget generation from community-asked questions and answers, and (3) nugget-level support, retrieving documents using a fusion of retrieval techniques and hybrid architectures. We use FreshStack to build five datasets on fast-growing, recent, and niche topics to ensure the tasks are sufficiently challenging. On FreshStack, existing retrieval models, when applied out-of-the-box, significantly underperform oracle approaches on all five topics, denoting plenty of headroom to improve IR quality. In addition, we identify cases where rerankers do not improve first-stage retrieval accuracy (two out of five topics) and oracle context helps an LLM generator generate a high-quality RAG answer. We hope FreshStack will facilitate future work toward constructing realistic, scalable, and uncontaminated IR and RAG evaluation benchmarks.
1 Introduction
Retrieval-augmented generation (RAG) is a popular technique to enhance traditional information retrieval (IR) capabilities with language model generation. RAG systems use large language models (LLMs) to generate long-form responses [20, 34, 23, 5], grounded in the information available from retrieved documents [31, 34, 18, 41]. Despite its wide usage, evaluating RAG remains incredibly challenging. Existing IR and RAG benchmarks are not well-suited for evaluation, as these are outdated and highly limited. In particular, we observe three major issues in existing benchmarks:
  • Lack of realistic, open-ended questions: Existing datasets contain purely extractive short answers (e.g., Natural Questions [33], TriviaQA [28]) or crowd-sourced questions (e.g., HotPotQA [83]). A limited number of datasets capture “natural” human-asked questions, i.e., MS MARCO [46] or Natural Questions [33], but unfortunately, brief and straightforward questions are inserted into a search box, failing to represent the complex questions that real users might pose to RAG systems.
@@ 1 Introduction
  • Static and unspecialized: After sourcing questions and answers, a benchmark becomes at the risk of (1) contamination, if current LLMs are trained on the same set of documents or questions, (2) overfitting, when systems inevitably saturate by repeated internal or external leaderboarding (e.g., BEIR [68]), and (3) staleness, when questions or answers are not refreshed and become outdated.
@@
Figure 1: A data instance from LangChain generated with FreshStack. The question and answer pair is sourced from Stack Overflow. The pair is provided to GPT-4o to generate nuggets, highlighting necessary facts in the answer. Next, code snippets and technical documents from multiple GitHub repositories (e.g., Jupyter Notebook) are chunked, processed, and pooled for each question. Finally, each pooled document chunk is judged with GPT-4o for binary relevance (either yes or no) at a nugget-level, i.e., whether the document factually supports the information present in each nugget.
@@
A realistic benchmark should measure model generalization on niche domains, and continue to update. Additionally, it must capture the complexity of human-generated queries—such as multi-hop reasoning [83], code understanding [27], or specialized terminology—rather than relying on artificially easy questions. This drives the robustness of systems in answering questions in evolving public libraries or private code bases [86, 27], a company’s internal forum [55], or technical troubleshooting [64, 53, 8].
@@
In our work, we introduce FreshStack, a holistic framework for constructing realistic datasets on niche and challenging domains, seeking to avoid contamination due to (perpetual) recency. Using FreshStack, we construct an evaluation benchmark on five niche topics sourced from community-asked questions and answers on Stack Overflow and a corpus containing code snippets and technical documents from public GitHub repositories. The framework contains three major steps:  (1) Automatic corpus collection (Section 3.2) with technical documents chunked and sourced from several GitHub repositories. (2) Nugget generation (Section 3.3) with GPT-4o using community-asked questions and answers in Stack Overflow. (3) Nugget-level support (Section 3.5) with GPT-4o on document chunks, retrieved from a fusion of retrieval techniques and hybrid architectures.
We investigate three research questions in our work to provide insights on FreshStack: (RQ1) How to construct challenging evaluation datasets from real user-asked questions? (RQ2) How do LLMs act as an assessor for nugget generation on community-asked questions & answers and nugget-level support with retrieved documents? (RQ3) How do state-of-the-art retrieval models, rerankers, and LLM generators perform on IR and RAG evaluation benchmarks generated with FreshStack?
@@
We calibrate the automatic stage in FreshStack with GPT-4o using a machine learning (ML) expert, assessing the quality of nugget generation and nugget-level support for one of the topics (LangChain). Our results show that GPT-4o-generated nuggets capture crucial information required to answer the question, and GPT-4o precisely labels support at a nugget level. For pooling, we compare oracle (having access to the answer) and inference (relying only on the question) settings, finding that question decomposition and nugget generation outperform other techniques, respectively.
@@
Beyond pool construction, we evaluate retrieval and rerankers in the document retrieval setting using only the Stack Overflow question. Retrieval models drastically underperform oracle systems on all five topics, showing a high headroom for improvement. In addition, ensemble fusion outperforms individual models, indicating that diversity in models enhances retrieval, and rerankers provide clear benefits in some but not all topics. Finally, we evaluate answer generation, where the oracle context assists the LLM generator to provide a high-quality RAG answer. FreshStack is a general framework and can be applied to any domain of a similar structure. Overall, we hope the framework serves as a testbed for future work to develop challenging benchmarks for evaluating RAG systems.
2 Related Work
@@
Retrieval-augmented generation. RAG has been widely used to avoid “hallucinations” [87] seen with LLMs when handling knowledge-intensive tasks [29]. RAG reduces factually incorrect generated content, leading to adoption in various commercial systems, e.g., Bing Search or Google AI Overviews. Existing IR and RAG benchmarks are stale, evaluating on academic question answering datasets [18, 58, 56, 65, 81], or are not challenging, being constructed for RAG [82, 7, 37, 47, 65, 32]. A limited number of datasets refresh over time to avoid LLM decontamination [30, 74, 62, 79], however, these contain easy and unrealistic questions. In contrast, FreshStack generates niche and challenging datasets, which can refresh over time and are not constructed specifically for RAG.
@@
Code-based benchmarks. Neural code generation [43] requires LLMs to generate code from scratch for generic programming questions. One popular benchmark is SWE-Bench [27], which evaluates whether LLMs can generate code changes for GitHub pull requests (PRs) in popular public repositories. Similarly, CodeSearchNet [22], COIR [36], LiveCodeBench [25], and CodeRAG-Bench [77] focus on the evaluation of high-level programming problems on popular public repositories. In contrast, in FreshStack, we focus on assisting developers, from a novice to a domain expert, by providing real-time answers on fast-growing and recent topics such as LangChain (introduced in 2023) by referencing technical documentation in GitHub repositories.
@@
Stack Overflow datasets. FreshStack is not the first dataset to use Stack Overflow for retrieval. However, the evaluation setting of retrieving canonical documents from GitHub repositories remains under-explored. Existing datasets such as CQADupstack [21], LoTTE [60], and Stack Overflow-QA [36] follow a different task, to retrieve the relevant answer snippet given the question asked by a real user on Stack Overflow. The closest setting similar to FreshStack is found in Neural Code Search [35], which incorporates public documentation and code snippet examples from GitHub as the corpus to answer questions asked by real users on popular programming topics such as Android.
@@
Figure 2: The FreshStack framework: (1) Stack Overflow questions and answers are sourced for recent and niche topics. (2) GitHub repository documents are collected and chunked to form the corpus (for each topic). (3) Nuggets or atomic facts within the question and answer are generated with GPT-4o. (4) Ensemble techniques and models retrieve documents, which construct our document judgment pools. (5) GPT-4o evaluates support for every document-nugget pair as a binary judgment.
3 The FreshStack Framework
@@
The framework involves five major stages to construct an evaluation dataset (as highlighted in Figure 2). FreshStack includes three major design choices:
  1. 1.
@@
    A general framework that can be extended to different domains without manual effort.
  2. 2.
    Adding recent and niche topics actively discussed in computer programmer communities such as Stack Overflow.
  3. 3.
@@
Stack Overflow is an online question answering platform for computer programmers. Users ask questions about a particular topic and provide a description (often a code snippet with the error message) with Stack Overflow tags. Questions and answers are also tagged by topics, allowing for easy retrieval of topic-wise questions.
3.1 Stack Overflow Topic and Question Selection
@@
For topic selection, we target niche and recent topics introduced on Stack Overflow from 2023 onward (frequency in Figure 3), containing a minimum of 50 posts. We sort all topics using the overall number of posts and curate five topics starting from the highest (LangChain) to the lowest frequency (Yolo v7 & v8), covering different domains and sufficiently different from each other. Each topic represents a unique domain, e.g., Computer Vision (CV) or Machine Learning (ML).
@@
Questions & Answers. We extract relevant posts and answers from the Stack Overflow XML data archive (dated October 2024). [^1] The Stack Overflow XML data archive (CC BY-SA license) is updated once every quarter: https://meta.stackexchange.com/questions/401324/announcing-a-change-to-the-data-dump-process We scan the archive to pick all the relevant posts as questions containing the required tag and further filter answer posts to the questions. Finally, we keep questions with an accepted answer, prioritizing precision over quantity in retaining questions with only high-quality answers accepted and upvoted in the Stack Overflow community.
3.2 Automatic Corpus Collection with GitHub Technical Documents
@@
Answering a higher percentage of questions requires a robust set of corpora from multiple sources. For instance, addressing issues in LangChain may require ChromaDB GitHub documentation to resolve errors related to its usage. In our work, we build a different document corpus per topic by combining multiple repositories as sources (we list each repository per topic in Table 7).
@@
Stack Overflow Tags. We analyze each tag frequency from Stack Overflow to select a relevant GitHub repository to be included in the document corpus. This involves identifying top-k co-occurring tags, where k is the threshold balancing question coverage with indexing costs. Some tags are generic, such as Python, whereas others are specific, such as LangChainJS. Filtering the tags to keep only a subset of repositories does not degrade the dataset quality. We manually verify each GitHub repository for each tag, with plans to automate this procedure in the future.
@@
Chunking & Indexing. We clone the latest branch of the GitHub repository in a local workspace and parse files as a tree structure. Each file (either a text document or code snippet) is chunked into small sections containing a maximum of 2048 tokens, skipping non-text formats. [^2] We skip indexing images, videos, .bin, .csv, and audio files or unrecognized file formats. The GitHub filepath serves as the document identifier, with additional chunk details encoded in the identifier. Finally, we combine all chunks into a single corpus, prefixing all document identifiers with the repository name to identify the common files in each repository separately (e.g., LICENSE or requirements.txt).
3.3 Nuggetization or Nugget Generation
@@
A nugget is a core concept or an atomic fact essential in a system’s response. The term nugget was informally referred to as SCU (summary content units) as clauses appearing in model summarization [45] and later formalized as “information nugget” for evaluating long-form answers [73, 38, 39, 49]. Nuggetization refers to constructing or generating nuggets from information-dense text. The procedure decomposes a verbose answer into key atomic facts or essential components, aiding evaluation. More recently, with the onset of RAG, nugget-based evaluation has renewed interest with LLMs for factual accuracy assessment in long answers [3, 44, 16, 50, 51, 52].
@@
Nuggetization. We automatically generate nuggets from Stack Overflow question-answer pairs using GPT-4o [48], avoiding the cumbersome procedure of manual nugget construction [50]. LLM-based nugget generation has been explored in the TREC 2024 RAG track [^3] TREC 2024 RAG track: https://trec-rag.github.io/ [50, 51, 52] and in multiple works [14, 16]. Separately, we experimented with prompting techniques and found that grading notes style prompts [42] provided parseable and high-quality nuggets in our experiments. An example of GPT-4o-generated nuggets for a question in LangChain is shown in Table 8.
3.4 Retrieval: Oracle & Inference Setting